Assume that Hart corporation's comparative balance sheet reported these amounts:

December 31

2017 2016

Plant and Equipment..................................... $625,000 $609,000

Less: Accumulated Depreciation................... $155,000 $135,000

Net Plant and Equipment............................... $470,000 $474,000

1. Assume that on January 1, 2018, Hart sold one-tenth of its plant and equipment for 56,000 cash. Journalize this transaction for Hart
Answer #1

Gross Plant & Equipment on 2017 = $625000

Accumulated Depreciation on 2017 = $155000

On January 1, 2018:

Cost of Plant & Equipment sold = 1/ 10th * Gross Plant & Equipment

Cost of Plant & Equipment sold = 1/10 * $625000 = $62500

Accumulated Depreciation on sold Equipment = 1 / 10th * Accumulated Depreciation

Accumulated Depreciation on sold Equipment = 1 / 10 * $155000 = $15500

We will now calculate the net book value of the Plant & Equipment sold:

Net book value of the Plant & Equipment sold = Cost of Plant & equipment sold - Accumulated Depreciation of the Equipment sold

Net book value of the Plant & Equipment sold = $62500 - $15500 = $47000

Now, we will calculate the gain or loss on sale as per below:

Gain (or loss) = Sale - Net book value

Gain (or loss) = $56000 - $47000 = $9000 (gain)

Required journal entry is:

Date Description Debit Credit
Cash 56000
Accumulated Depreciation 15500
Gain on sale 9000
Plant & Equipment 62500